# VDGD: Mitigating LVLM Hallucinations in Cognitive Prompts by Bridging the Visual Perception Gap

### v1

@@ H.4 Additional Details on Algorithm 1
    Next, if the Base Rank is more than 0 and the hallucination can be found in the top-k retrieved elements from the IT dataset the model was aligned using, we attribute it to be an IT hallucination [8]. To retrieve top-k image-instruction pairs that are closest to the prompt image, we employ CLIP [25] similarity. Specifically, we calculate the similarity between the image embedding of the image associated with the input prompt and the text embedding of the description/response associated with the IT instances. Next, for every prompt, we retrieve the top-25 IT instances and search the hallucinated phrase returned by the judge (in descriptions/responses) using string matching.
@@
H.5 Elbow Method for Logit Analysis
@@
H.6 Examples of Rephrased Prompts
To determine whether LVLM hallucinations stem from the limitations of the base LLM or from alignment issues, we ask GPT-4 to rewrite the instructions in prompts in such a way that the base LLM can answer the question without any image input as described in section 5.2. We share a few examples (figures: 30, 31, 32, 33, 34, 35, 36, 37) of rephrased prompts in the existing datasets below:
@@ H.6 Examples of Rephrased Prompts
Methodology VP VP+Info-Seek VP+Info-Seek+Reasoning VP+Reasoning
LLaVA-v1 1.62 2.60 1.67 1.70
LLaVA-v1.5 2.59 3.18 1.50 1.76
LLaVA-v1.6 3.46 3.68 1.67 2.15
mPLUG-Owl2 2.56 3.23 1.67 1.94
InternLM-X 3.75 3.90 2.33 2.56
CogVLM 3.00 3.37 1.25 2.08
@@
Dataset Model k Variance Range Avg.
AMBER LLaVA-v1 4.19 0.05 0.31 0.23
LLaVA-v1.5 3.91 0.07 0.39 0.29
LLaVA-v1.6 3.78 0.08 0.43 0.32
mPLUG-Owl2 3.65 0.06 0.36 0.28
InternLM-X 3.56 0.09 0.44 0.36
CogVLM 3.16 0.91 0.47 0.38
MMMU LLaVA-v1 3.68 0.08 0.39 0.27
LLaVA-v1.5 3.53 0.09 0.41 0.23
LLaVA-v1.6 3.40 0.09 0.42 0.24
mPLUG-Owl2 3.32 0.10 0.44 0.26
InternLM-X 3.26 0.14 0.46 0.28
CogVLM 3.16 0.19 0.47 0.25
@@
Dataset Model Helpfulness Clarity Factuality Depth Engagement
VaLLu LLaVA-v1.5-t 3.05 4.02 2.83 2.70 3.85
mPLUG-Owl2-t 3.70 4.15 3.14 3.11 3.24
InternLM-X-t 4.28 4.58 3.66 3.62 3.57
MMMU LLaVA-v1.5-t 1.97 3.25 1.69 1.87 3.25
mPLUG-Owl2-t 2.48 3.05 1.63 2.22 2.64
InternLM-X-t 3.51 3.93 2.43 2.98 3.01
MathVista LLaVA-v1.5-t 2.53 3.58 2.21 2.10 3.44
mPLUG-Owl2-t 3.56 4.05 3.04 2.52 2.79
InternLM-X-t 4.39 4.65 3.99 2.91 2.95
Oven LLaVA-v1.5-t 3.38 4.48 2.95 2.32 3.35
mPLUG-Owl2-t 3.62 4.56 2.99 2.47 3.37
InternLM-X-t 3.80 4.58 3.76 2.89 3.52
Dataset Model Helpfulness Clarity Factuality Depth Engagement
AMBER LLaVA-v1 3.79 4.46 3.2 3.27 0.96
LLaVA-1.5 4.29 4.73 3.91 3.44 1.26
LLaVA-1.6 4.30 4.78 3.85 4.29 1.27
mPLUG-Owl2 4.02 4.58 3.48 3.27 0.71
InternLM-X 4.59 4.86 4.37 4.00 1.40
CogVLM 4.68 4.93 4.49 3.96 1.28
GPT-4-Turbo 4.89 4.96 4.79 4.64 1.50
OCR LLaVA-v1 1.11 2.09 1.11 1.14 0.47
LLaVA-1.5 1.07 2.24 1.17 1.15 0.48
LLaVA-1.6 2.51 2.92 2.63 1.99 1.11
mPLUG-Owl2 1.21 1.15 1.16 1.10 0.64
InternLM-X 1.11 1.12 1.12 1.03 0.67
CogVLM 3.12 3.04 2.68 2.53 1.47
GPT-4-Turbo 4.15 4.02 3.78 3.98 1.67
MMMU LLaVA-v1 1.77 2.73 1.50 1.65 0.46
LLaVA-1.5 2.45 3.34 2.00 2.23 0.59
LLaVA-1.6 3.54 3.98 2.95 3.39 0.99
mPLUG-Owl2 2.94 3.64 2.56 2.33 0.82
InternLM-X 3.60 4.10 3.24 3.15 1.05
CogVLM 4.03 4.38 3.57 3.43 1.03
GPT-4-Turbo 4.86 4.92 4.73 4.56 1.15
MathVista LLaVA-v1 1.95 2.84 1.95 1.85 0.76
LLaVA-1.5 2.57 3.47 2.23 2.36 0.98
LLaVA-1.6 3.68 4.07 3.17 3.48 1.12
mPLUG-Owl2 3.14 3.85 2.78 2.46 1.03
InternLM-X 3.91 4.32 3.45 3.35 1.23
CogVLM 4.20 4.52 3.78 3.66 1.24
GPT-4-Turbo 4.54 4.97 4.12 4.10 1.35
MMC LLaVA-v1 1.81 2.90 1.51 1.75 0.56
LLaVA-1.5 2.66 3.49 2.21 2.42 0.71
LLaVA-1.6 4.09 4.32 3.50 3.90 1.09
mPLUG-Owl2 3.63 4.15 3.32 2.60 0.97
InternLM-X 3.72 4.17 3.28 3.25 1.10
CogVLM 4.27 4.58 3.91 3.92 1.32
GPT-4-Turbo 4.81 4.89 4.57 4.49 1.12
Model Methodology Helpfulness Clarity Factuality Depth Engagement
LLaVA-v1 Vanilla-greedy 2.61 3.69 1.95 1.92 1.25
Vanilla-sampling 2.64 3.73 1.86 1.94 1.27
VDD 1.70 2.55 1.47 1.81 1.59
OPERA 2.60 3.75 2.04 1.89 1.30
Woodpecker 2.68 3.81 2.01 1.95 1.28
LRV 2.64 3.78 1.98 1.94 1.36
LURE 2.60 3.54 2.03 1.89 1.40
VDGD (ours) 2.38 3.39 2.16 2.46 1.34
LLaVA-1.5 Vanilla-greedy 2.64 3.72 2.03 2.03 1.25
Vanilla-sampling 2.57 3.65 2.01 1.92 1.22
VDD 1.94 2.99 1.55 1.65 1.05
OPERA 2.73 3.74 2.05 2.04 1.22
Woodpecker 2.74 3.87 2.05 2.03 1.29
LRV 2.76 3.82 2.10 2.03 1.28
LURE 2.72 3.76 2.03 2.05 1.27
VDGD (ours) 2.97 3.91 2.64 2.27 1.54
LLaVA-1.6 Vanilla-greedy 3.08 3.81 2.63 2.67 1.43
Vanilla-sampling 3.10 3.85 2.64 2.63 1.40
VDD 2.01 2.96 1.80 2.10 1.20
OPERA 3.10 4.01 2.62 2.60 1.42
Woodpecker 3.16 4.00 2.67 2.58 1.44
LRV 3.15 3.98 2.65 2.62 1.39
LURE 3.11 4.02 2.64 2.64 1.48
VDGD (ours) 3.51 4.04 3.16 2.87 1.58
mPLUG-Owl2 Vanilla-greedy 2.73 3.64 2.20 1.96 1.28
Vanilla-sampling 2.72 3.60 2.18 1.91 1.24
VDD 2.16 2.95 1.80 1.85 1.26
OPERA 2.92 3.96 2.26 2.08 1.34
Woodpecker 2.89 3.90 2.23 2.02 1.24
LRV 2.90 3.94 2.19 2.05 1.28
LURE 2.92 3.89 2.24 2.06 1.32
VDGD (ours) 3.14 3.98 2.72 2.22 1.50
InternLM-X Vanilla-greedy 3.11 3.98 2.66 2.34 1.47
Vanilla-sampling 3.13 4.04 2.70 2.41 1.56
VDD 2.56 3.26 2.32 2.19 1.41
OPERA 3.14 4.20 2.65 2.32 1.46
Woodpecker 3.13 4.12 2.60 2.28 1.41
LRV 3.06 4.19 2.59 2.31 1.45
LURE 3.12 4.10 2.64 2.40 1.49
VDGD (ours) 3.57 4.37 3.45 2.65 1.68
CogVLM Vanilla-greedy 3.25 4.19 2.82 2.43 1.55
Vanilla-sampling 3.26 4.21 2.83 2.40 1.55
VDD 2.80 3.62 2.63 2.10 1.58
OPERA 3.40 4.27 2.85 2.46 1.57
Woodpecker 3.44 4.30 2.91 2.45 1.59
LRV 3.42 4.26 2.88 2.42 1.54
LURE 3.39 4.20 2.78 2.41 1.52
VDGD (ours) 3.33 4.15 3.01 2.51 1.60
GPT-4-Turbo - 4.02 4.44 3.63 2.94 1.59
@@
Model Methodology Helpfulness Clarity Factuality Depth Engagement
LLaVA-v1 Vanilla-greedy 1.81 2.79 1.36 1.51 1.00
Vanilla-sampling 1.82 2.82 1.38 1.56 1.06
VDD 1.70 2.67 1.22 1.40 0.95
OPERA 1.88 2.87 1.45 1.61 1.07
Woodpecker 1.89 2.88 1.48 1.60 1.06
LRV 1.83 2.85 1.40 1.57 1.08
LURE 1.91 2.90 1.52 1.63 1.09
VDGD (ours) 2.09 2.99 1.64 1.74 1.14
LLaVA-1.5 Vanilla-greedy 1.86 2.90 1.47 1.70 1.09
Vanilla-sampling 1.83 2.88 1.42 1.67 1.08
VDD 1.70 2.52 1.34 1.58 0.97
OPERA 2.18 3.08 1.62 1.75 1.22
Woodpecker 2.15 2.93 1.57 1.63 1.18
LRV 2.16 2.88 1.60 1.67 1.16
LURE 2.03 2.87 1.46 1.71 1.15
VDGD (ours) 2.30 3.17 1.85 1.86 1.20
LLaVA-1.6 Vanilla-greedy 1.85 2.67 1.54 1.80 0.97
Vanilla-sampling 1.80 2.63 1.50 1.78 0.96
VDD 1.68 2.57 1.36 1.68 0.94
OPERA 1.86 2.70 1.57 1.84 0.99
Woodpecker 1.90 2.74 1.63 1.89 1.04
LRV 1.88 2.72 1.61 1.85 1.02
LURE 1.92 2.73 1.65 1.86 1.03
VDGD (ours) 2.11 2.89 1.81 1.92 1.05
mPLUG-Owl2 Vanilla-greedy 2.24 3.12 1.58 1.65 1.16
Vanilla-sampling 2.25 3.12 1.59 1.66 1.14
VDD 2.10 2.98 1.42 1.56 1.04
OPERA 2.35 3.23 1.67 1.74 1.18
Woodpecker 2.28 3.17 1.61 1.69 1.15
LRV 2.34 3.24 1.66 1.72 1.14
LURE 2.32 3.21 1.63 1.67 1.19
VDGD (ours) 2.48 3.36 1.82 1.78 1.23
InternLM-X Vanilla-greedy 2.95 3.71 2.14 2.39 1.63
Vanilla-sampling 2.97 3.76 2.17 2.45 1.64
VDD 2.76 3.54 2.03 2.26 1.57
OPERA 3.09 3.86 2.20 2.43 1.66
Woodpecker 3.07 3.83 2.19 2.46 1.68
LRV 3.11 3.88 2.24 2.49 1.70
LURE 3.10 3.89 2.22 2.48 1.71
VDGD (ours) 3.21 4.11 2.57 2.78 1.79
CogVLM Vanilla-greedy 2.42 3.34 1.69 1.83 1.22
Vanilla-sampling 2.44 3.37 1.68 1.84 1.23
VDD 2.22 3.15 1.53 1.68 1.17
OPERA 2.56 3.47 1.76 1.88 1.27
Woodpecker 2.57 3.47 1.78 1.90 1.26
LRV 2.58 3.49 1.80 1.92 1.25
LURE 2.54 3.42 1.75 1.85 1.22
VDGD (ours) 2.65 3.56 1.92 1.97 1.29
GPT-4-Turbo - 4.15 4.53 3.66 2.99 1.95
@@
Table 8: Performance comparison of various LVLMs on MMMU benchmark. This is an extension of results shown in Figure 7.
Model Methodology Helpfulness Clarity Factuality Depth Engagement
LLaVA-v1 Vanilla-greedy 3.74 4.54 2.23 2.28 1.42
Vanilla-sampling 3.72 4.56 2.24 2.27 1.43
VDD 3.54 4.32 2.05 2.07 1.36
OPERA 3.80 4.63 2.29 2.35 1.50
Woodpecker 3.78 4.60 2.26 2.30 1.47
LRV 3.76 4.64 2.30 2.38 1.49
LURE 3.74 4.59 2.25 2.28 1.46
VDGD (ours) 3.92 4.78 2.52 2.42 1.54
LLaVA-1.5 Vanilla-greedy 3.65 4.38 2.36 2.01 1.32
Vanilla-sampling 3.68 4.39 2.38 2.04 1.36
VDD 3.38 4.09 2.35 1.87 1.10
OPERA 3.74 4.46 2.26 1.93 1.32
Woodpecker 3.46 4.53 2.34 1.97 1.38
LRV 3.53 4.42 2.39 1.96 1.27
LURE 3.33 4.41 2.37 1.91 1.29
VDGD (ours) 3.84 4.61 2.56 2.15 1.45
LLaVA-1.6 Vanilla-greedy 3.72 4.38 2.50 2.34 1.50
Vanilla-sampling 3.69 4.32 2.45 2.28 1.48
VDD 3.54 4.11 2.34 2.18 1.35
OPERA 3.80 4.47 2.59 2.46 1.57
Woodpecker 3.78 4.42 2.56 2.43 1.53
LRV 3.75 4.40 2.54 2.42 1.52
LURE 3.79 4.41 2.52 2.41 1.53
VDGD (ours) 3.90 4.53 2.79 2.51 1.58
mPLUG-Owl2 Vanilla-greedy 3.79 4.43 2.48 2.08 1.33
Vanilla-sampling 3.80 4.45 2.51 2.12 1.34
VDD 3.62 4.27 2.35 1.98 1.24
OPERA 3.85 4.52 2.57 2.20 1.40
Woodpecker 3.84 4.56 2.60 2.24 1.42
LRV 3.82 4.54 2.59 2.22 1.41
LURE 3.86 4.58 2.61 2.25 1.45
VDGD (ours) 3.99 4.65 2.76 2.36 1.52
InternLM-X Vanilla-greedy 4.02 4.69 2.84 2.35 1.55
Vanilla-sampling 4.01 4.68 2.83 2.37 1.54
VDD 3.82 4.37 2.66 2.10 1.44
OPERA 4.13 4.76 2.92 2.42 1.59
Woodpecker 4.07 4.72 2.88 2.40 1.58
LRV 4.10 4.73 2.89 2.40 1.56
LURE 4.14 4.75 2.91 2.43 1.60
VDGD (ours) 4.28 4.89 3.12 2.54 1.66
CogVLM Vanilla-greedy 3.72 4.39 2.76 2.63 1.44
Vanilla-sampling 3.73 4.38 2.74 2.61 1.42
VDD 3.58 4.27 2.64 2.50 1.37
OPERA 3.76 4.45 2.82 2.69 1.47
Woodpecker 3.75 4.44 2.79 2.64 1.45
LRV 3.80 4.51 2.86 2.70 1.49
LURE 3.82 4.53 2.88 2.73 1.52
VDGD (ours) 3.94 4.64 2.99 2.80 1.60
GPT-4-Turbo - 4.05 4.75 3.17 2.24 1.61

### v2

@@ Appendix I Model Details
    Next, if the Base Rank is more than 0 and the hallucination can be found in the top-k retrieved elements from the IT dataset the model was aligned using, we attribute it to be an IT hallucination (Ghosh et al., 2024). To retrieve top-k image-instruction pairs that are closest to the prompt image, we employ CLIP (Radford et al., 2021) similarity. Specifically, we calculate the similarity between the image embedding of the image associated with the input prompt and the text embedding of the description/response associated with the IT instances. Next, for every prompt, we retrieve the top-25 IT instances and search the hallucinated phrase returned by the judge (in descriptions/responses) using string matching. We show how varying values of k change the final counts for our hallucination categories in Appendix K.5.
@@
K.5 Results for various values of k for IT Hallucinations
Fig. 33 shows that even for higher and lower values of k, our hypothesis and claims stay unchanged as changing the value of k does not affect the overall frequency distribution of categories.
[…]
[width=0.5]chart-2.pdf
Figure 33: Frequency comparison of hallucination categories for top-k=10,25,40
K.6 Elbow Method for Logit Analysis
@@
K.7 Examples of Rephrased Prompts
To determine whether LVLM hallucinations stem from the limitations of the base LLM or from alignment issues, we ask GPT-4 to rewrite the instructions in prompts in such a way that the base LLM can answer the question without any image input as described in section 4.2. We share a few examples (figures: 34, 35, 36, 37, 38, 39, 40, 41) of rephrased prompts in the existing datasets below:
[…]
[width=]vallu_rephrased_img_1.pdf
@@
K.8 Compute Infrastructure
Appendix L Additional Results
L.1 Comparison of various LVLMs on rephrased prompts
[…]
! Dataset Model Helpfulness Clarity Factuality Depth Engagement VaLLu LLaVA-v1.5-t 3.05 4.02 2.83 2.70 3.85 mPLUG-Owl2-t 3.70 4.15 3.14 3.11 3.24 InternLM-X-t 4.28 4.58 3.66 3.62 3.57 MMMU LLaVA-v1.5-t 1.97 3.25 1.69 1.87 3.25 mPLUG-Owl2-t 2.48 3.05 1.63 2.22 2.64 InternLM-X-t 3.51 3.93 2.43 2.98 3.01 MathVista LLaVA-v1.5-t 2.53 3.58 2.21 2.10 3.44 mPLUG-Owl2-t 3.56 4.05 3.04 2.52 2.79 InternLM-X-t 4.39 4.65 3.99 2.91 2.95 Oven LLaVA-v1.5-t 3.38 4.48 2.95 2.32 3.35 mPLUG-Owl2-t 3.62 4.56 2.99 2.47 3.37 InternLM-X-t 3.80 4.58 3.76 2.89 3.52
L.2 Comparison of various LVLMs on VaLLu benchmark with no image input.
Table 5 provides comparison of various LVLMs on VaLLU benchmark when prompted with no image but image descriptions are provided.
[…]
! Methodology LLaVA-v1 LLaVA-v1.5 LLaVA-v1.6 mPLUG-Owl2 InternLM-X CogVLM Vanilla-greedy 1.67 1.78 2.07 1.76 2.02 2.23 VCD 1.31 1.29 1.40 1.42 2.04 2.21 OPERA 1.65 1.43 2.29 1.91 2.30 2.44 Woodpecker 1.77 1.85 2.29 2.02 2.41 2.60
Table 5: Comparison of LVLMs performance when prompted without an image but image descriptions are provided on VaLLu benchmark. The values correspond to the "Factuality" metric.
L.3 Comparison of various LVLMs on image description
L.4 Comparison of LVLMs on different capabilities on the proposed VaLLu benchmark
Table 7 provides a comparison of LVLMs on different capabilities on the proposed VaLLu benchmark.
[…]
! Dataset Model Helpfulness Clarity Factuality Depth Engagement AMBER LLaVA-v1 3.79 4.46 3.2 3.27 0.96 LLaVA-1.5 4.29 4.73 3.91 3.44 1.26 LLaVA-1.6 4.30 4.78 3.85 4.29 1.27 mPLUG-Owl2 4.02 4.58 3.48 3.27 0.71 InternLM-X 4.59 4.86 4.37 4.00 1.40 CogVLM 4.68 4.93 4.49 3.96 1.28 GPT-4-Turbo 4.89 4.96 4.79 4.64 1.50 OCR LLaVA-v1 1.11 2.09 1.11 1.14 0.47 LLaVA-1.5 1.07 2.24 1.17 1.15 0.48 LLaVA-1.6 2.51 2.92 2.63 1.99 1.11 mPLUG-Owl2 1.21 1.15 1.16 1.10 0.64 InternLM-X 1.11 1.12 1.12 1.03 0.67 CogVLM 3.12 3.04 2.68 2.53 1.47 GPT-4-Turbo 4.15 4.02 3.78 3.98 1.67 MMMU LLaVA-v1 1.77 2.73 1.50 1.65 0.46 LLaVA-1.5 2.45 3.34 2.00 2.23 0.59 LLaVA-1.6 3.54 3.98 2.95 3.39 0.99 mPLUG-Owl2 2.94 3.64 2.56 2.33 0.82 InternLM-X 3.60 4.10 3.24 3.15 1.05 CogVLM 4.03 4.38 3.57 3.43 1.03 GPT-4-Turbo 4.86 4.92 4.73 4.56 1.15 MathVista LLaVA-v1 1.95 2.84 1.95 1.85 0.76 LLaVA-1.5 2.57 3.47 2.23 2.36 0.98 LLaVA-1.6 3.68 4.07 3.17 3.48 1.12 mPLUG-Owl2 3.14 3.85 2.78 2.46 1.03 InternLM-X 3.91 4.32 3.45 3.35 1.23 CogVLM 4.20 4.52 3.78 3.66 1.24 GPT-4-Turbo 4.54 4.97 4.12 4.10 1.35 MMC LLaVA-v1 1.81 2.90 1.51 1.75 0.56 LLaVA-1.5 2.66 3.49 2.21 2.42 0.71 LLaVA-1.6 4.09 4.32 3.50 3.90 1.09 mPLUG-Owl2 3.63 4.15 3.32 2.60 0.97 InternLM-X 3.72 4.17 3.28 3.25 1.10 CogVLM 4.27 4.58 3.91 3.92 1.32 GPT-4-Turbo 4.81 4.89 4.57 4.49 1.12
[…]
0.95! Methodology VP VP+Info-Seek VP+Info-Seek+Reasoning VP+Reasoning LLaVA-v1 1.62 2.60 1.67 1.70 LLaVA-v1.5 2.59 3.18 1.50 1.76 LLaVA-v1.6 3.46 3.68 1.67 2.15 mPLUG-Owl2 2.56 3.23 1.67 1.94 InternLM-X 3.75 3.90 2.33 2.56 CogVLM 3.00 3.37 1.25 2.08
@@
L.5 Comparison of post-truncation probability statistics on AMBER and MMMU benchmarks
[…]
0.7! Dataset Model k Variance Range Avg. AMBER LLaVA-v1 4.19 0.05 0.31 0.23 LLaVA-v1.5 3.91 0.07 0.39 0.29 LLaVA-v1.6 3.78 0.08 0.43 0.32 mPLUG-Owl2 3.65 0.06 0.36 0.28 InternLM-X 3.56 0.09 0.44 0.36 CogVLM 3.16 0.91 0.47 0.38 MMMU LLaVA-v1 3.68 0.08 0.39 0.27 LLaVA-v1.5 3.53 0.09 0.41 0.23 LLaVA-v1.6 3.40 0.09 0.42 0.24 mPLUG-Owl2 3.32 0.10 0.44 0.26 InternLM-X 3.26 0.14 0.46 0.28 CogVLM 3.16 0.19 0.47 0.25
@@
L.6 Comparison of VaLLu with Mitigation Baselines
L.7 Comparison of MMMU with Mitigation Baselines
L.8 Comparison of MathVista with Mitigation Baselines
L.9 Comparison of Oven with Mitigation Baselines
[…]
! Model Methodology Helpfulness Clarity Factuality Depth Engagement LLaVA-v1 Vanilla-greedy 2.61 3.69 1.95 1.92 1.25 Vanilla-sampling 2.64 3.73 1.86 1.94 1.27 VDD 1.70 2.55 1.47 1.81 1.59 OPERA 2.60 3.75 2.04 1.89 1.30 Woodpecker 2.68 3.81 2.01 1.95 1.28 LRV 2.64 3.78 1.98 1.94 1.36 LURE 2.60 3.54 2.03 1.89 1.40 VDGD (ours) 2.38 3.39 2.16 2.46 1.34 LLaVA-1.5 Vanilla-greedy 2.64 3.72 2.03 2.03 1.25 Vanilla-sampling 2.57 3.65 2.01 1.92 1.22 VDD 1.94 2.99 1.55 1.65 1.05 OPERA 2.73 3.74 2.05 2.04 1.22 Woodpecker 2.74 3.87 2.05 2.03 1.29 LRV 2.76 3.82 2.10 2.03 1.28 LURE 2.72 3.76 2.03 2.05 1.27 VDGD (ours) 2.97 3.91 2.64 2.27 1.54 LLaVA-1.6 Vanilla-greedy 3.08 3.81 2.63 2.67 1.43 Vanilla-sampling 3.10 3.85 2.64 2.63 1.40 VDD 2.01 2.96 1.80 2.10 1.20 OPERA 3.10 4.01 2.62 2.60 1.42 Woodpecker 3.16 4.00 2.67 2.58 1.44 LRV 3.15 3.98 2.65 2.62 1.39 LURE 3.11 4.02 2.64 2.64 1.48 VDGD (ours) 3.51 4.04 3.16 2.87 1.58 mPLUG-Owl2 Vanilla-greedy 2.73 3.64 2.20 1.96 1.28 Vanilla-sampling 2.72 3.60 2.18 1.91 1.24 VDD 2.16 2.95 1.80 1.85 1.26 OPERA 2.92 3.96 2.26 2.08 1.34 Woodpecker 2.89 3.90 2.23 2.02 1.24 LRV 2.90 3.94 2.19 2.05 1.28 LURE 2.92 3.89 2.24 2.06 1.32 VDGD (ours) 3.14 3.98 2.72 2.22 1.50 InternLM-X Vanilla-greedy 3.11 3.98 2.66 2.34 1.47 Vanilla-sampling 3.13 4.04 2.70 2.41 1.56 VDD 2.56 3.26 2.32 2.19 1.41 OPERA 3.14 4.20 2.65 2.32 1.46 Woodpecker 3.13 4.12 2.60 2.28 1.41 LRV 3.06 4.19 2.59 2.31 1.45 LURE 3.12 4.10 2.64 2.40 1.49 VDGD (ours) 3.57 4.37 3.45 2.65 1.68 CogVLM Vanilla-greedy 3.25 4.19 2.82 2.43 1.55 Vanilla-sampling 3.26 4.21 2.83 2.40 1.55 VDD 2.80 3.62 2.63 2.10 1.58 OPERA 3.40 4.27 2.85 2.46 1.57 Woodpecker 3.44 4.30 2.91 2.45 1.59 LRV 3.42 4.26 2.88 2.42 1.54 LURE 3.39 4.20 2.78 2.41 1.52 VDGD (ours) 3.33 4.15 3.01 2.51 1.60 GPT-4-Turbo - 4.02 4.44 3.63 2.94 1.59
@@
[…]
! Model Methodology Helpfulness Clarity Factuality Depth Engagement LLaVA-v1 Vanilla-greedy 1.81 2.79 1.36 1.51 1.00 Vanilla-sampling 1.82 2.82 1.38 1.56 1.06 VDD 1.70 2.67 1.22 1.40 0.95 OPERA 1.88 2.87 1.45 1.61 1.07 Woodpecker 1.89 2.88 1.48 1.60 1.06 LRV 1.83 2.85 1.40 1.57 1.08 LURE 1.91 2.90 1.52 1.63 1.09 VDGD (ours) 2.09 2.99 1.64 1.74 1.14 LLaVA-1.5 Vanilla-greedy 1.86 2.90 1.47 1.70 1.09 Vanilla-sampling 1.83 2.88 1.42 1.67 1.08 VDD 1.70 2.52 1.34 1.58 0.97 OPERA 2.18 3.08 1.62 1.75 1.22 Woodpecker 2.15 2.93 1.57 1.63 1.18 LRV 2.16 2.88 1.60 1.67 1.16 LURE 2.03 2.87 1.46 1.71 1.15 VDGD (ours) 2.30 3.17 1.85 1.86 1.20 LLaVA-1.6 Vanilla-greedy 1.85 2.67 1.54 1.80 0.97 Vanilla-sampling 1.80 2.63 1.50 1.78 0.96 VDD 1.68 2.57 1.36 1.68 0.94 OPERA 1.86 2.70 1.57 1.84 0.99 Woodpecker 1.90 2.74 1.63 1.89 1.04 LRV 1.88 2.72 1.61 1.85 1.02 LURE 1.92 2.73 1.65 1.86 1.03 VDGD (ours) 2.11 2.89 1.81 1.92 1.05 mPLUG-Owl2 Vanilla-greedy 2.24 3.12 1.58 1.65 1.16 Vanilla-sampling 2.25 3.12 1.59 1.66 1.14 VDD 2.10 2.98 1.42 1.56 1.04 OPERA 2.35 3.23 1.67 1.74 1.18 Woodpecker 2.28 3.17 1.61 1.69 1.15 LRV 2.34 3.24 1.66 1.72 1.14 LURE 2.32 3.21 1.63 1.67 1.19 VDGD (ours) 2.48 3.36 1.82 1.78 1.23 InternLM-X Vanilla-greedy 2.95 3.71 2.14 2.39 1.63 Vanilla-sampling 2.97 3.76 2.17 2.45 1.64 VDD 2.76 3.54 2.03 2.26 1.57 OPERA 3.09 3.86 2.20 2.43 1.66 Woodpecker 3.07 3.83 2.19 2.46 1.68 LRV 3.11 3.88 2.24 2.49 1.70 LURE 3.10 3.89 2.22 2.48 1.71 VDGD (ours) 3.21 4.11 2.57 2.78 1.79 CogVLM Vanilla-greedy 2.42 3.34 1.69 1.83 1.22 Vanilla-sampling 2.44 3.37 1.68 1.84 1.23 VDD 2.22 3.15 1.53 1.68 1.17 OPERA 2.56 3.47 1.76 1.88 1.27 Woodpecker 2.57 3.47 1.78 1.90 1.26 LRV 2.58 3.49 1.80 1.92 1.25 LURE 2.54 3.42 1.75 1.85 1.22 VDGD (ours) 2.65 3.56 1.92 1.97 1.29 GPT-4-Turbo - 4.15 4.53 3.66 2.99 1.95
@@
Table 10: Performance comparison of various LVLMs on MMMU benchmark (only questions tagged with open-ended generation). This is an extension of results shown in Figure 7.
[…]
! Model Methodology Helpfulness Clarity Factuality Depth Engagement LLaVA-v1 Vanilla-greedy 3.74 4.54 2.23 2.28 1.42 Vanilla-sampling 3.72 4.56 2.24 2.27 1.43 VDD 3.54 4.32 2.05 2.07 1.36 OPERA 3.80 4.63 2.29 2.35 1.50 Woodpecker 3.78 4.60 2.26 2.30 1.47 LRV 3.76 4.64 2.30 2.38 1.49 LURE 3.74 4.59 2.25 2.28 1.46 VDGD (ours) 3.92 4.78 2.52 2.42 1.54 LLaVA-1.5 Vanilla-greedy 3.65 4.38 2.36 2.01 1.32 Vanilla-sampling 3.68 4.39 2.38 2.04 1.36 VDD 3.38 4.09 2.35 1.87 1.10 OPERA 3.74 4.46 2.26 1.93 1.32 Woodpecker 3.46 4.53 2.34 1.97 1.38 LRV 3.53 4.42 2.39 1.96 1.27 LURE 3.33 4.41 2.37 1.91 1.29 VDGD (ours) 3.84 4.61 2.56 2.15 1.45 LLaVA-1.6 Vanilla-greedy 3.72 4.38 2.50 2.34 1.50 Vanilla-sampling 3.69 4.32 2.45 2.28 1.48 VDD 3.54 4.11 2.34 2.18 1.35 OPERA 3.80 4.47 2.59 2.46 1.57 Woodpecker 3.78 4.42 2.56 2.43 1.53 LRV 3.75 4.40 2.54 2.42 1.52 LURE 3.79 4.41 2.52 2.41 1.53 VDGD (ours) 3.90 4.53 2.79 2.51 1.58 mPLUG-Owl2 Vanilla-greedy 3.79 4.43 2.48 2.08 1.33 Vanilla-sampling 3.80 4.45 2.51 2.12 1.34 VDD 3.62 4.27 2.35 1.98 1.24 OPERA 3.85 4.52 2.57 2.20 1.40 Woodpecker 3.84 4.56 2.60 2.24 1.42 LRV 3.82 4.54 2.59 2.22 1.41 LURE 3.86 4.58 2.61 2.25 1.45 VDGD (ours) 3.99 4.65 2.76 2.36 1.52 InternLM-X Vanilla-greedy 4.02 4.69 2.84 2.35 1.55 Vanilla-sampling 4.01 4.68 2.83 2.37 1.54 VDD 3.82 4.37 2.66 2.10 1.44 OPERA 4.13 4.76 2.92 2.42 1.59 Woodpecker 4.07 4.72 2.88 2.40 1.58 LRV 4.10 4.73 2.89 2.40 1.56 LURE 4.14 4.75 2.91 2.43 1.60 VDGD (ours) 4.28 4.89 3.12 2.54 1.66 CogVLM Vanilla-greedy 3.72 4.39 2.76 2.63 1.44 Vanilla-sampling 3.73 4.38 2.74 2.61 1.42 VDD 3.58 4.27 2.64 2.50 1.37 OPERA 3.76 4.45 2.82 2.69 1.47 Woodpecker 3.75 4.44 2.79 2.64 1.45 LRV 3.80 4.51 2.86 2.70 1.49 LURE 3.82 4.53 2.88 2.73 1.52 VDGD (ours) 3.94 4.64 2.99 2.80 1.60 GPT-4-Turbo - 4.05 4.75 3.17 2.24 1.61